OpenStudy (anonymous):

I'm really bad with word problems, please help! What quantity of a 50% acid solution must be mixed with a 35% solution to produce 300 mL of a 45% solution?

OpenStudy (anonymous):

x = quantity of 50 % acid solution y = quanity of 35% acid solution x + y = 300 ml 0.5 * x + 0.35 * y = 0.45 * 300 ml 0.5 x + 0.35 y = 135 ml of acid 50 x + 35 y = 13500 10 x + 7 y = 2700 x + y = 300 => 10x + 10y = 3000 10x + 7 y = 2700 3y =3000 - 2700 3y = 300 y = 100 x + 100 = 300 => x = 200